OpenStudy (anonymous):

solve y''+y'-6y=12(e^(3t))+12e^(-2t)

OpenStudy (anonymous):

Homogeneous solution: \[y''+y'-6y=0\] yields the characteristic equation \[r^2+r-6=0~~\Rightarrow~~r=-3,2\] So homogeneous part is \(y_c=C_1e^{-3t}+C_2e^{2t}\). Non-homogeneous solution: \[y''+y'-6y=12e^{3t}+12e^{-2t}\] Use the method of undetermined coefficients. Suppose \(y_p=Ae^{3t}+Be^{-2t}\) is a solution, so you have \[\begin{cases} y_p=Ae^{3t}+Be^{-2t}\\ y_p'=3Ae^{3t}-2Be^{-2t}\\ y_p'=9Ae^{3t}+4Be^{-2t} \end{cases}\] Substitute into the original equation: \[\begin{align*}\left(9Ae^{3t}+4Be^{-2t}\right)+\left(3Ae^{3t}-2Be^{-2t}\right)-6\left(Ae^{3t}+Be^{-2t}\right)&=12e^{3t}+12e^{-2t}\\ 6Ae^{3t}-4Be^{-2t}&=12e^{3t}+12e^{-2t} \end{align*}\] This tells you that \(A=2\) and \(B=-3\), and so your non-homogeneous part is \(y_p=2e^{3t}-3e^{-2t}\). Your final solution would be the sum of the non/homogeneous parts, or \[y=y_c+y_p\]